What is the brightness (total power output) of the Aldebaran star
(we admit that the luminous intensity of Aldebaran is isotropic)?

we know that the parallaxe is 0,049’’
and the bolometric brightness is 3,4x10^-8 W/m2

thank you :)
Answer #1

We know that,

Apparent brightness = Actual brightness/4πd²

Here, d is the distance from the star

I parsec = 3.086 × 10^13 km

49 parsec = 1.512 × 10^15 km

Rearranging the above formulae

Actual brightness = apparent brightness × 4πd²

On putting the values,

Actual brightness = 34×10^-8 × 4 × π × (1.512 × 10^15)²

Actual brightness = 646.012 × 10^22

Actual brightness = 6.46 × 10^24 W/m²